OpenStudy (candy13106):

4-2y-y=12 solve the equation

OpenStudy (anonymous):

solve for y =.=

OpenStudy (candy13106):

yea i kinda need help

OpenStudy (cj49):

4-3y=12 -3y=12-4 -3y=8 y=-8/3

OpenStudy (candy13106):

i got y=-2 2/3

OpenStudy (cj49):

-2y-y=-3y

OpenStudy (cj49):

can u show ur work

OpenStudy (candy13106):

yea i know

OpenStudy (candy13106):